If Sally can paint a house in 4 hours, and John can paint the same house in 6 hours, how long will it take for both of them to paint the house together?
Answer by edjones(7311)

Sally can do 1/4 of the job in 1 hr.
John can do 1/6 of the job in 1 hr.
1/6 + 1/4
=2/12 + 3/12 = 5/12 of the job in 1 hr together so they can do the entire job in 12/5 hrs.
12/5=2 2/5 hrs or 2 hrs 24 min.
.
Ed